OpenStudy (anonymous):

find an equation for the line that passes through the point (2,-5) and a) is parallel to the x-axis b) is parallel to the y-axis c) is parallel to the line 2x-4y=3

OpenStudy (hoblos):

parallel to the x-axis ---> horizontal line of slope=0 so y=-5 parallel to the y-axis ---> vertical line of undefined slope so x=2 parallel lines have same slope 2x-4y=3 4y=2x-3 y=(1/2)x -3/4 ------> slope=1/2 so y+5=1/2(x-2) y+5 = (1/2)x -1 y=(1/2)x -6

OpenStudy (anonymous):

ok , thank you , i got it

OpenStudy (hoblos):

yw :)

OpenStudy (anonymous):

b) y = -5 parallel to x-axis means y is constant, since we have (2,-5) on the line y = -5 is the line c) x = 2 parallel to y-axis means x is constant, since we have (2,-5) on the line x = 2 is the line d) 2x - 4y = 3 4y = 2x - 3 y = 1/2x - 3/4 You see the slope is 1/2 i.e m For our line y = mx + b to be parallel to y = 1/2x - 3/4 both should have the same slope i.e 1/2 y = 1/2x + b Substitute x = 2 and y =-5 -5 = 1/2 * 2 + b -5 = 1 + b b = -6 So our line is y = 1/2x - 6

OpenStudy (anonymous):

thank you so much